MICROWAVE SPECTRUM AND STRUCTURE OF THE 2,6-DIFLUOROPYRIDINE-CO_{2} COMPLEX

\underline{CHRISTOPHER T. DEWBERRY}, Department of Chemistry, University of Minnesota, 207 Pleasant St., SE, Minneapolis, MN 55455; JESSICA L. MUELLER, MARK D. MARSHALL, HELEN O. LEUNG, Department of Chemistry, Amherst College, P.O. Box 5000, Amherst, MA 01002-5000; KENNETH R. LEOPOLD, Department of Chemistry, University of Minnesota, 207 Pleasant St., SE, Minneapolis, MN 55455.

The microwave spectrum of the weakly bound complex formed from 2,6-difluoropyridine and CO_{2} has been obtained using the Amherst broadband microwave spectrometer. Preliminary analysis indicates a structure in which the nitrogen of the 2,6-difluoropyridine is directed toward the carbon of the CO_{2}. However, unlike the related complex pyridine-CO_{2}, which is planar, the CO_{2} is rotated out of the plane of the heterocycle. The nitrogen-carbon van der Waals distance is 2.916 Å, significantly larger than the 2.7977(64) Å value previously reported for pyridine-CO_{2}. Ab initio calculations at the MP2/6-311++G(2d,2p) level support a nitrogen bound geometry, but indicate an equilibrium structure in which the CO_{2} lies off the C_{2} axis of the 2,6-difluoropyridine. One dimensional discrete variable representation (DVR) calculations using an ab initio potential restricted to rotation of CO_{2} about the C_{2} axis of the difluoropyridine indicate that the complex is in the high-barrier limit to internal rotation and that no splittings due to internal rotation will be observed in the spectrum.